OpenStudy (anonymous):

Solve for x. e^x e^(x+1)=1

OpenStudy (anonymous):

\[e ^{x}e ^{x+1}=1\] \[e ^{x+x+1}=1,e ^{2x+1}=1,2x+1=\ln 1=0,x=-\frac{ 1 }{ 2 }\]